# PePNet: A Periodicity-Perceived Workload Prediction Network Supporting Rare Occurrence of Heavy Workload

### v2

@@ II-B2 Periodicity-Fusing Module
To filter out the random noise in the periodic information, PePNet firstly uses an auto-encoder [22]. To solve the problem of variable periodicity for different machines and the problem of local periodicity violation, PePNet uses an attention mechanism to evaluate the reliability of periodic information. The final prediction $y$ is given in Eq.13-14.
$y_{period}^{{}^{\prime}}=Autoencoder(y_{period})$ (13)
$y=Attention(X^{enc}_{short},(\hat{y},y_{period}^{{}^{\prime}}),(\hat{y},y_{period}^{{}^{\prime}}))$ (14)
II-C Achilles’ Heel Loss Function
Heavy load rarely occurs, which leads to poor performance of heavy load prediction due to data imbalance. To solve this problem, PePNet deploys the Achilles’ Heel Loss Function. As analyzed below, the Achilles’ Heel Loss Function actually forces the loss to descend along the gradient directions of some of the largest prediction errors. When iteratively training the model, the Achilles’ Heel Loss Function optimizes the most under-fitting part in one step. In the next step, the original largest prediction error diminishes and the most under-fitting part shifts to another place. In this way, the Achilles’ Heel Loss Function optimizes each part of the sequence iteratively and effectively improves the accuracy of heavy load, when keeping high predicting accuracy of overall workload. The Achilles’ Heel Loss Function is given in Eq.15, where $T$ equals $\{t_{1},\ldots,t_{1+J}\}$, which are the time slots of a forecasting sequence. $y_{t}$ denotes the predicted workload at time $t$ (the prediction length for one sample should be at least 2), $Y_{t}$ denotes the ground-truth workload at time $t$, and $\gamma$ is a hyperparameter.
@@ II-C Achilles’ Heel Loss Function
The Achilles’ Heel Loss Function actually distributes more weight to the gradient for higher prediction error when back propagating. This aspect is obvious by deriving Eq.15, as shown in Eq.16, where $\alpha_{t}=\frac{exp{\frac{(y_{t}(\mathcal{P})-Y_{t})^{2}}{\gamma}}}{\sum_{t=1}^{T}\exp{\frac{(y_{t}(\mathcal{P})-Y_{t})^{2}}{\gamma}}}$, $\mathcal{P}$ stands for the parameters of PePNet. $\alpha_{t}$ is actually a normalized weight, whose value depends on the prediction error at time slot $t$. $\gamma$ is a scale factor. When $\gamma$ is large, the difference in prediction errors across time slots is narrowed and the weights are relatively uniform for each time slot. When the $\gamma$ is small, the difference in prediction errors across time slots is amplified and the weights become polarized. To visually demonstrate this point, we plot the value of $\alpha_{t}$ for different combinations of $\gamma$ and prediction error $y_{t}-Y_{t}$ in Fig. 3(h). In Fig. 3(h), each row shows the value of $\alpha_{t}$ for a set of prediction error $\{y_{t}-Y_{t}|0\leq t\leq 9\}$, when specifying the $\gamma$. As Fig. 3(h) shown, for a specific $\gamma$, the larger the prediction error is, the bigger the $\alpha_{t}$ is. Besides, for a specific prediction error in a specific error set, the smaller the $\gamma$ is, the bigger the $\alpha_{t}$ is. When $\gamma$ is set to 0.1, $\alpha_{t}$ for the largest prediction error in this row approximates 1.
@@
It is worth noticing that when the $\gamma$ is indefinitely close to 0, the Achilles’ Heel Loss Function becomes a simple formation, which is given in the following theorem. The theorem implies that when $\gamma$ indefinitely approximates 0, only the gradient for maximum prediction error is assigned a weight of 1, while others are assigned a weight of 0.
Theorem 2. As the $\gamma$ is infinitely close to 0, the gradient of the loss function infinitely approximates Eq.17. The $t^{*}=\mathop{\arg\max}_{t}(y_{t}-Y_{t})^{2}$ and the $g(t^{*},\mathcal{P})=(y[t^{*}]-Y[t^{*}])^{2}$.
$\frac{\partial l(Y,y)}{\partial\mathcal{P}}=\frac{\partial g(t^{*},\mathcal{P})}{\partial\mathcal{P}}$ (17)
Proof of Theorem 2. The derivation of $l(Y,y)$ is shown in Eq.18. Then, both the numerator and dominator on the right-hand side are divided by $\exp(\frac{g(t^{*},\mathcal{P})}{\gamma})$ simultaneously and the first step of Eq.19 is obtained. The $g(t^{*},\mathcal{P})$ is surely bigger than $g(t,\mathcal{P}),\forall t\neq t^{*}$, as the definition of $t^{*}$ suggests. Thus, $g(t,\mathcal{P})-g(t^{*},\mathcal{P})<0,\forall t\neq t^{*}$. When the $\gamma(\gamma>0)$ infinitely approaches zero, $exp(\frac{g(t,\mathcal{P})-g(t^{*},\mathcal{P})}{\gamma})$ will also infinitely approaches zero. In this way, the second step of Eq.19 is obtained.
$=\frac{\partial g(t^{*},\mathcal{P})}{\partial\mathcal{P}}$
@@ III Experiment
In this section, we conduct extensive experiments to validate the following findings:
  • PePNet improves the accuracy of heavy-workload prediction as well as the overall prediction accuracy compared with the state-of-the-art methods.
  • PePNet only introduces slight time overhead for training and inference.
  • PePNet is insensitive to hyperparameters and has high robustness robustness.
  • Each mechanism in PePNet is proven to play an important role by using ablation experiments.
  • Due to its adaptability for periodicity, when the time series is aperiodic, PePNet still works well.
TABLE III: Basic hyperparameters of PePNet.
@@
Number of layers of LSTM 2
Input length of Alibaba2018 for short-term information 50
Input length of Alibaba2018 for long-term information 100
Input length of Dinda’s dataset for short-term information 30
Input length of Dinda’s dataset for long-term information 80
Input length of SMD for short-term information 50
Input length of SMD for long-term information 100
Hidden layer size of LSTM 80
Predicted length 2
Init downsampling step size 20
Number of layers of auto-encoder 1
Number of layers of auto-decoder 1
Output size of auto-encoder 1
Sliding window length for downsampling 100
@@ III-A Experiment Setup
Hyperparameters. We summarize the most important hyperparameters of PePNet in Tab. III.
Baseline Methods. We compare PePNet with the state-of-the-art time series predicting methods and popular workload predicting models and the variants of PePNet.
  • Autoformer (Autof) [23]: Autoformer is a novel decomposition architecture with an Auto-Correlation mechanism. It breaks the pre-processing convention of series decomposition and renovates it as a basic inner block of deep models. Further, Autoformer also has an Auto-Correlation mechanism based on the series periodicity.
  • LSTM with attention mechanism (LSTMa) [10]: The approach extracts the sequential and contextual features of the historical workload data by an encoder network and integrates an attention mechanism into the decoder network.
  • Informer (Inf) [11]: The method uses a ProbSparse attention mechanism and a halving cascading layer to extract the input information and use a generative style decoder to predict. Informer is one of the most recognized methods of time series prediction in recent years.
  • L-PAW [5]: The method integrates top-sparse auto-encoder (TSA) and gated recurrent unit (GRU) block into RNN to achieve the adaptive and accurate prediction for highly-variable workloads. L-PAW is a model specifically designed for workload prediction and is also widely used.
  • LSTM [20]: A classic time series processing model, which uses the hidden state to store long sequences of information.
  • GRU [18]: A variant of the LSTM that modifies the forget gate structure in LSTM to make the model simpler.
  • Reformer (Ref) [24]: Reformer introduces two techniques to improve the efficiency of Transformers. Firstly, it replaces dot-product attention with one that uses locality-sensitive hashing. Secondly, it uses reversible residual layers instead of standard residuals to reduce the memory consumption of the model. It is a widely recognized method in time series prediction.
  • Fedformer (Fedf) [25]: Fedformer combines Transformer with a seasonal-trend decomposition method, in which the decomposition method captures the global profile of time series while Transformers capture more detailed structures. Fedformer is an effective and novel method in time series prediction.
@@
  • Variants of PePNet: We implement the backbone-encoder and backbone-decoder as transformer and LSTM respectively and denote them as PePNet-I and PePNet-II. Without specification, we use PePNet to denote PePNet-II in the following. Moreover, we also introduce some variants for ablation study. We use PePNet- to denote PePNet-II that removes the Periodicity-Perceived Mechanism. We use PePNet to denote the PePNet-II that uses MSE as a loss function. We use PePNet to denote the PePNet-II that both removes the Periodicity-Perceived Mechanism and does not use the Achilles’ Heel loss function.
TABLE IV: The overall and heavy-workload prediction accuracy on three datasets. The MSE in the table are multiplied by 1000.
LSTM LSTMa GRU Inf L-PAW Ref Autof Fedf PePNet- PePNet PePNet PePNet-I PePNet-II
MAPE 0.093 0.117 0.108 0.100 0.102 0.104 0.211 0.110 0.089 0.087 0.116 0.047 0.076
Overall MSE 5.515 8.233 9.516 4.865 10.960 8.950 37.636 9.211 5.773 6.146 6.847 4.420 3.697
Dinda’s MAE 0.048 0.071 0.070 0.041 0.078 0.053 0.149 0.057 0.046 0.049 0.059 0.054 0.029
Dataset MAPE 0.051 0.079 0.094 0.030 0.046 0.040 0.079 0.032 0.035 0.043 0.058 0.065 0.018
Heavy-Workload MSE 12.260 17.925 24.191 6.837 15.853 9.202 23.817 8.729 5.635 7.325 13.274 8.760 6.124
MAE 0.068 0.104 0.123 0.040 0.058 0.051 0.097 0.041 0.046 0.057 0.076 0.083 0.025
MAPE 0.179 0.238 0.287 0.237 0.296 0.278 0.488 0.290 0.207 0.191 0.161 0.076 0.154
Overall MSE 3.434 3.627 3.889 3.282 3.909 4.950 8.021 2.350 3.144 2.974 3.101 0.552 1.757
SMD MAE 0.019 0.023 0.025 0.022 0.025 0.028 0.046 0.023 0.020 0.019 0.018 0.015 0.014
Dataset MAPE 0.150 0.163 0.154 0.166 0.154 0.316 0.367 0.157 0.138 0.129 0.158 0.204 0.112
Heavy-Workload MSE 10.437 10.622 11.945 11.024 11.912 17.587 27.465 5.531 10.941 9.751 10.399 4.138 6.729
MAE 0.037 0.045 0.046 0.041 0.046 0.068 0.095 0.034 0.042 0.038 0.040 0.055 0.033
MAPE 0.039 0.095 0.049 0.058 0.157 0.086 0.037 0.025 0.050 0.027 0.126 0.017 0.016
Overall MSE 0.375 0.489 0.462 0.408 12.721 0.880 1.173 0.510 0.359 0.341 0.975 0.410 0.323
Alibaba MAE 0.014 0.017 0.015 0.015 0.058 0.023 0.026 0.016 0.014 0.013 0.024 0.014 0.013
Memory MAPE 0.023 0.023 0.027 0.027 0.039 0.022 0.032 0.030 0.019 0.019 0.030 0.019 0.019
Heavy-Workload MSE 0.706 0.673 0.860 0.919 2.420 0.695 1.418 1.182 0.566 0.542 1.114 0.598 0.509
MAE 0.021 0.020 0.023 0.024 0.032 0.019 0.029 0.027 0.018 0.018 0.028 0.018 0.017
MAPE 0.173 0.215 0.181 0.165 0.272 0.190 0.258 0.158 0.139 0.158 0.150 0.153 0.142
Overall MSE 4.647 4.996 4.753 4.809 12.638 6.097 9.954 6.204 4.786 4.812 4.784 6.039 4.758
Alibaba MAE 0.049 0.053 0.051 0.050 0.090 0.060 0.073 0.056 0.050 0.051 0.050 0.054 0.049
CPU MAPE 0.195 0.237 0.194 0.193 0.343 0.247 0.381 0.252 0.178 0.177 0.182 0.176 0.143
Heavy-Workload MSE 16.767 17.548 16.228 17.742 29.493 18.848 32.271 22.226 16.594 16.321 16.405 14.090 11.634
MAE 0.095 0.100 0.096 0.100 0.150 0.105 0.145 0.115 0.097 0.095 0.095 0.087 0.082
@@
(e) Heavy-workload MAE of PePNet on Alibaba2018-CPU
(f) Overall MAE of PePNet on Alibaba2018-Memory
(g) Heavy-workload MAE of PePNet on Alibaba2018-Memory
(h) Overall MAE of PePNet on Dinda’s dataset
(i) Heavy-workload MAE of PePNet on Dinda’s dataset
@@
Figure 4: (a) The value of the y-axis is the original workload divided by the maximum workload of its machine. The wider the width of the violin in the figure is, the more workloads are distributed near this value. (b) The figure shows the average consuming time of one forward propagation and backpropagation for one sample. The labels on the histogram are 1000 times of the actual value. Besides, we use the shorthand of Informer, Autoformer and Reformer in the xticks. (c) The figure shows the overall and heavy-workload MAE of PePNet for different $\gamma$ on all the three datasets. In the legend, we use dataset name* to represent the heavy-workload MAE on the specific dataset, while using dataset name to represent the overall MAE on the dataset. (d)-(i) These figures show overall and heavy-workload MAE of PePNet on different datasets with different combination of the length of $X_{in}$ and $p$.
Datasets. We perform experiments on three public datasets: Alibaba’s cluster trace v2018444https://github.com/alibaba/clusterdata, Dinda’s dataset [12] and Server Machine Dataset (SMD) [26]. These datasets are collected by different organizations with different system configurations. Their use can help to verify the generalization of PePNet. We plot the data distribution of different datasets in Fig. 4(a). It is worth noting that different datasets represent different data distributions: long-tailed distribution (Dinda’s dataset), centralized distribution (memory usage of Alibaba2018 and SMD), and relatively even data distribution (CPU usage of Alibaba2018).
  • Dinda’s dataset (long-tailed distribution): Dinda’s dataset is collected from two groups of machines. The first is the Alpha cluster at the Pittsburgh Supercomputing Center (PSC). The second one includes computing servers (Mojave, Sahara), a testbed (manchester1-8), and desktop workstations at Carnegie Mellon University. The workload in Dinda’s dataset shows significant periodicity and sometimes there are large peaks far above average (long-tailed distribution).
  • Alibaba2018-Memory (centralized distribution): Alibaba’s cluster trace is sampled from one of Alibaba’s production clusters, which includes about 4000 machines’ workload in 8 days. Memory usage of some machines in this dataset shows relatively significant periodicity and others are aperiodic. Memory usage is generally stable and is distributed around a certain value (centralized distribution).
  • Alibaba2018-CPU (relatively even distribution): The CPU usage in Alibaba2018 shows more significant periodicity on the heavy workload than the light load. Besides, it is highly variable and is distributed relatively even.
  • SMD dataset (centralized distribution with rare extreme values): SMD is a new 5-week-long dataset, which is collected from a top 10 Internet company. SMD is made up of data from 28 different machines. It consists of 38 features, including CPU utilization, memory utilization, disk I/O, etc. The workload in this dataset shows high periodicity. Besides, most of the workload of this dataset is concentrated around 10-20, while others are extremely big.
We use the first 1,000,000 lines in Alibaba’s dataset and axp0 in Dinda’s dataset. The values in Alibaba2018 are utilization percentage and we uniformly scale it to range [0, 1].
Evaluation metrics. We use three metrics to evaluate PePNet’s performance: Mean Squared Error (MSE), Mean Absolute Error (MAE), and Mean Absolute Percentage Error (MAPE). These metrics are some of the most recognized and widely used ones in time series prediction. There are many influential works using these metrics, such as [11], [5], [23]. Besides, these metrics not only measure the relative error (MAPE) but also the absolute error (MSE, MAE).

### v3

@@ II-B2 Periodicity-Fusing Module
To filter out random noise in the periodic information, PePNet firstly uses an auto-encoder [27]. To solve the problem of variable periodicity for different machines and the problem of local periodicity violation, PePNet uses an attention mechanism to evaluate the reliability of periodic information. The final prediction $y$ is given as:
$y_{p}^{{}^{\prime}}=\operatorname{Autoencoder}(y_{p})$ (8)
$y=\operatorname{Attention}(\bar{X},(\hat{y},y_{p}^{{}^{\prime}}),(\hat{y},y_{p}^{{}^{\prime}}))$ (9)
II-C Achilles’ Heel Loss Function (AHLF)
Heavy loads occur infrequently, causing severe data imbalance and poor prediction performance. To address this issue, PePNet adopts the AHLF, which focus on the part with largest prediction error. During training, it prioritizes the most under-fitted regions. As these errors are reduced, the focus adaptively shifts to other under-fitted parts. This iterative process effectively improves heavy-load prediction accuracy while maintaining high overall forecasting performance. AHLF is given as:
@@ II-C Achilles’ Heel Loss Function (AHLF)
where $T=\{t_{1},\ldots,t_{1+J}\}$, which are the time slots of a forecasting sequence. $y_{t}$ denotes the predicted workload at time $t$ (the prediction length for one sample should be at least 2), $Y_{t}$ denotes the ground-truth workload at time $t$, and $\gamma$ is a scale factor.
AHLF actually distributes more weight to the gradient for higher prediction error when back propagating. This aspect is obvious by taking first deriving of Eq. (10), i.e:
@@
where $\alpha_{t}=\frac{exp{\frac{(y_{t}(\mathcal{P})-Y_{t})^{2}}{\gamma}}}{\sum_{t=1}^{T}\exp{\frac{(y_{t}(\mathcal{P})-Y_{t})^{2}}{\gamma}}}$, $\mathcal{P}$ stands for the parameters of PePNet. $\alpha_{t}$ is actually a normalized weight whose value depends on the prediction error at time slot $t$. $\gamma$ is a scale factor. When $\gamma$ is large, the difference in prediction errors across time slots is narrowed and the weights are relatively uniform for each time slot. When $\gamma$ is small, the difference in prediction errors across time slots is amplified and the weights become polarized. To visually demonstrate this point, we plot the value of $\alpha_{t}$ for different combinations of $\gamma$ and prediction error $y_{t}-Y_{t}$ in Fig. 3(d). In it, each row shows the value of $\alpha_{t}$ for a set of prediction error $\{y_{t}-Y_{t}|0\leq t\leq 9\}$, when specifying $\gamma$. For a specific $\gamma$, the larger the prediction error is, the bigger $\alpha_{t}$ is. Besides, for a specific prediction error in a specific error set, the smaller $\gamma$ is, the bigger $\alpha_{t}$ is. When $\gamma$ is set to 0.1, $\alpha_{t}$ for the largest prediction error in this row approximates 1.
It is worth noticing that when $\gamma$ is indefinitely close to 0, AHLF becomes a simple formation, as given in the following theorem.
Theorem 2. As $\gamma$ is infinitely close to 0, the gradient of the loss function infinitely approximates:
$\frac{\partial g(t^{*},\mathcal{P})}{\partial\mathcal{P}}$ (12)
where $t^{*}=\mathop{\arg\max}_{t}(y_{t}-Y_{t})^{2}$ and $g(t^{*},\mathcal{P})=(y[t^{*}]-Y[t^{*}])^{2}$.
Proof of Theorem 2 is given in Appendix. B.
The theorem implies that when $\gamma$ indefinitely approximates 0, only the gradient for maximum prediction error is assigned a weight of 1, while others are assigned a weight of 0.
@@ III Experiment
In this section, we conduct extensive experiments to validate the following findings: 1) PePNet improves the accuracy of heavy-workload prediction as well as the overall one compared with the state-of-the-art methods; 2) PePNet only introduces slight time overhead for training and inference over the existing methods; 3) PePNet is insensitive to hyperparameters and has high robustness; 4) Each mechanism in PePNet is proven to play an important role via ablation studies; and 5) Due to its adaptability for periodicity, when the time series is aperiodic, PePNet still works well.
TABLE I: Basic hyperparameters of PePNet.
@@ III-A Experiment Setup
Hyperparameters. We summarize the most important hyperparameters of PePNet in Tab. I.
@@
Baseline Methods. We compare PePNet with the state-of-the-art time series predicting methods and popular workload predicting models and the variants of PePNet : Autoformer (Autof) [35], LSTM with attention mechanism (LSTMa) [45], Informer (Inf) [42], L-PAW [8], LSTM [18], GRU [9], Reformer (Ref) [22], Fedformer (Fedf) [43], and Time-MoE (TMoE) [30]. Besides, we implement the backbone-encoder and backbone-decoder as transformer and LSTM respectively and denote them as PeP-T and PePNet. Moreover, we also introduce some variants for ablation study. We use PeP- to denote PePNet that removes the Periodicity-Perceived Mechanism. We use PeP to denote the PePNet that uses MSE as a loss function. We use PeP to denote the PePNet that both removes the Periodicity-Perceived Mechanism and does not use AHLF but uses MSE.
Datasets. We perform experiments on three datasets: Alibaba’s cluster trace v2018333https://github.com/alibaba/clusterdata, Dinda’s dataset collected from Pittsburgh Supercomputing Center [12] and Server Machine Dataset (SMD) [31] collected from one of global top-10 Internet companies. From different organizations with different system configurations, they can help us to verify the generalization of PePNet. We plot the data distribution of different datasets in Fig. 4(a). It is worth noting that different datasets represent different data distributions: long-tailed distribution (Dinda’s dataset), centralized distribution (memory usage of Alibaba2018 and SMD), and relatively even data distribution (CPU usage of Alibaba2018). We use the first 1,000,000 lines in Alibaba’s dataset and axp0 in Dinda’s dataset. The values in Alibaba2018 are utilization percentage and we uniformly scale it to range [0, 1].
Evaluation metrics. We use three metrics to evaluate performance of all the methods: MSE, Mean Absolute Error (MAE), and Mean Absolute Percentage Error (MAPE).
@@
Figure 4: (a) The value of the y-axis is the original workload divided by the maximum workload of its machine. The wider the width of the violin in the figure is, the more workloads are distributed near this value. (b) The figure shows the average consuming time of one forward propagation and backpropagation for one sample. The labels on the histogram are 1000 times of the actual value. Besides, we use the shorthand of Informer, Autoformer and Reformer in the xticks. (c) The figure shows the overall and heavy-workload MAE of PePNet for different $\gamma$ on all the three datasets. In the legend, we use dataset name* to represent the heavy-workload MAE on the specific dataset, while using dataset name to represent the overall MAE on the dataset. (d) The figure show overall MAE of PePNet with different combination of the length of $X_{in}$ and $p$.
TABLE II: The overall and heavy-workload prediction accuracy on three datasets. The MSE in the table are multiplied by 1000.
LSTMa Inf L-PAW Ref Autof Fedf TMoE PeP- PeP PeP PeP-T PePNet
MAPE 0.117 0.100 0.102 0.104 0.211 0.110 0.752 0.089 0.087 0.116 0.047 0.076
Overall MSE 8.233 4.865 10.960 8.950 37.636 9.211 5.171 5.773 6.146 6.847 4.420 3.697
Dinda’s MAE 0.071 0.041 0.078 0.053 0.149 0.057 0.061 0.046 0.049 0.059 0.054 0.029
Dataset MAPE 0.079 0.030 0.046 0.040 0.079 0.032 0.769 0.035 0.043 0.058 0.065 0.018
Heavy MSE 17.925 6.837 15.853 9.202 23.817 8.729 15.504 5.635 7.325 13.274 8.760 6.124
MAE 0.104 0.040 0.058 0.051 0.097 0.041 0.123 0.046 0.057 0.076 0.083 0.025
MAPE 0.238 0.237 0.296 0.278 0.488 0.290 0.951 0.207 0.191 0.161 0.076 0.154
Overall MSE 3.627 3.282 3.909 4.950 8.021 2.350 2.012 3.144 2.974 3.101 0.552 1.757
SMD MAE 0.023 0.022 0.025 0.028 0.046 0.023 0.040 0.020 0.019 0.018 0.015 0.014
Dataset MAPE 0.163 0.166 0.154 0.316 0.367 0.157 0.689 0.138 0.129 0.158 0.204 0.112
Heavy MSE 10.622 11.024 11.912 17.587 27.465 5.531 3.652 10.941 9.751 10.399 4.138 6.729
MAE 0.045 0.041 0.046 0.068 0.095 0.034 0.056 0.042 0.038 0.040 0.055 0.033
MAPE 0.095 0.058 0.157 0.086 0.037 0.025 0.135 0.050 0.027 0.126 0.017 0.016
Overall MSE 0.489 0.408 12.721 0.880 1.173 0.510 14.168 0.359 0.341 0.975 0.410 0.323
Alibaba MAE 0.017 0.015 0.058 0.023 0.026 0.016 0.056 0.014 0.013 0.024 0.014 0.013
Memory MAPE 0.023 0.027 0.039 0.022 0.032 0.030 0.069 0.019 0.019 0.030 0.019 0.019
Heavy MSE 0.673 0.919 2.420 0.695 1.418 1.182 8.624 0.566 0.542 1.114 0.598 0.509
MAE 0.020 0.024 0.032 0.019 0.029 0.027 0.065 0.018 0.018 0.028 0.018 0.017
MAPE 0.215 0.165 0.272 0.190 0.258 0.158 0.266 0.139 0.158 0.150 0.153 0.142
Overall MSE 4.996 4.809 12.638 6.097 9.954 6.204 7.033 4.786 4.812 4.784 6.039 4.758
Alibaba MAE 0.053 0.050 0.090 0.060 0.073 0.056 0.060 0.050 0.051 0.050 0.054 0.049
CPU MAPE 0.237 0.193 0.343 0.247 0.381 0.252 0.342 0.178 0.177 0.182 0.176 0.143
Heavy MSE 17.548 17.742 29.493 18.848 32.271 22.226 38.502 16.594 16.321 16.405 14.090 11.634
MAE 0.100 0.100 0.150 0.105 0.145 0.115 0.153 0.097 0.095 0.095 0.087 0.082